OpenStudy (anonymous):

two dear run toward each other . one deer has a mass of 20kg and runs west with a speed of 5 m/s, while the other has a mass of 30 kg and runs east with a speed of 4 m/s. what is the total momentum of the system made up of two deer? a. 20 kg.m/s east b.220kg.m/s west c.20 kg.m/s west d. 220 kg.m/s east

OpenStudy (anonymous):

Momentum of deer 1 is: 20kg x 5m/s West = 100 kgm/s West or -100kgm/s East Momentum of deer 2 is: 30kg x 4m/s East = 120 kgm/s East Total (net) momentum Momentum of deer 1 + Momentum of deer 2 taken in the same direction (i.e. East) -100+120 = 20 kgm/s East